OpenStudy (anonymous):

450 C of charge flows through a motor and 9000 J of energy are converted in the motor. 1800 J are dissipated in the cell. What is the emf?

OpenStudy (anonymous):

as u might know, W=qV. so in this case 9000+1800=450(emf or the V of eqn). ie emf=10800/450=60/25=2.4 volts